OpenStudy (anonymous):

Write an equation in point slope form and slop intercept. (-3,-3) and (2,7)

OpenStudy (anonymous):

slope = (7-(-3))/(2-(-3)) = 10/5=2 y-(-3)=(slope)(x-(-3)) y+3=2(x+3) y=2x+6-3 y=2x+3

OpenStudy (anonymous):

you need an explanation?

OpenStudy (anonymous):

yes plz smartbum is that the slope form or slope intercept

OpenStudy (anonymous):

Of what ol dude wrote above?

OpenStudy (anonymous):

point slope form is written in this form: y - y1= m(x-x1) and slope intercept form is written in this form : y= mx + b

OpenStudy (anonymous):

so y=2x + 3 is slope intercept

OpenStudy (anonymous):

and y+ 3= 2(x + 3) is point slope

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

no problem